OpenStudy (anonymous):

Which number is halfway between 1 half and 5 over 6?

OpenStudy (anonymous):

Which number is halfway between 1/2 and 5/6?

OpenStudy (anonymous):

1/2= 0.5 5/6=0.833333333 7/10=0.7 2/3= 0.666666666 27/40=0.675

OpenStudy (anonymous):

@Godlovesme